Kosar Manufacturing has collected the following information on one of its major products.

Most efficient production rate = 2,200 + X units per period

Production change costs = $30 per unit of change (from the most efficient production rate = 2,100 + X units/period)

Inventory costs = $7 per unit per period (on closing inventory balance)

Backorder costs = $15 per unit to carry demand into next period

X= 32

Period

Demand (units)

1

3,000 + X

2

2,300

3

1,800 + X

4

2,800

5

2,000 + X

6

2,300

Beginning inventory = “X*20” units

Calculate a level production schedule that yields X inventory at the end of period 6.

Total

Ending

Regular

Extra

Quarter

Demand

Production

Inventory

Production

Production

0

1

2

3

4

5

6

What is the total cost of this solution? What is the total inventory cost of this solution? (25 points)

Extra production

Inventory

Backorder

Costs

Total cost
Answer #1

a. Since X = 32, efficient production rate = 2,200 + 32 = 2,232 units and beginning inventory = 32 * 20 = 640 units.

We shall aim to produce maximum number of units in a given period irrespective of the demand.

Quarter Demand Production Ending inventory Regular Production Extra Production
640
1 3032 2392 0 2232 160
2 2300 2232 68 2232
3 1832 2232 400 2232
4 2800 2400 0 2232 168
5 2032 2232 0 2232
6 2300 2332 32 2232 100
Average 2382.67 2303.33 83.33 2232 142.67

b.

Quarter Demand Production Ending inventory (A) Regular Production Extra Production (B) Extra Production@30/unit (B * 30) Inventory @$7/unit (A * 7) Backorder @15/unit (B*15)
640 4480
1 3032 2392 0 2232 160 4800 0 2400
2 2300 2232 68 2232 476
3 1832 2232 400 2232 2800
4 2800 2400 0 2232 168 5040 2520
5 2032 2232 0 2232
6 2300 2332 32 2232 100 3000 224 1500
Average 2382.67 2303.33 83.33 2232 142.67 12840 7980 6420